What is the formal charge on atom in carbonate and nitric acid?

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
The charge assigned to an atom in a molecule is a formal charge.
Formal charge(FC)is given by the formula,
FC=V−N−B/2
Where,
V=Number of valence electrons
N=Number of non-bonding electrons
B=Total number of electrons shared in covalent bonds.
The formal charge ofNinNO2−is
FC=5−2−6/2=0
For the LHS oxygen theFCis-
FC=6−6−2/2=−1
For the RHS oxygen the FC is-
FC=6−4−4/2=0
The carbonate ion[CO3​2−]has an overall charge of -2. In the carbonate ion, the carbon atom is bonded with a double bond to an oxygen atom, and with single bonds to two oxygen atoms.
Formal charge on oxygen bonded with the double bond =6−4−1/2(4)=0
Formal charge on oxygen atoms bonded with single bond =6−6−1/2(2)=−1
InHNO3​
The nitrogen atom carries a charge of+1, and one oxygen atom carries a charge of−1.